OpenStudy (anonymous):

50th term is 238, 93rd term is 539. state the general term

hartnn (hartnn):

of AP?

OpenStudy (anonymous):

what?

hartnn (hartnn):

of Arithmetic sequence ??

hartnn (hartnn):

general term of what?

OpenStudy (anonymous):

yes arithmetic

hartnn (hartnn):

the general 'n' term is \[T_n=a+((n-1)d)\] so T50=238=a+49d ----->(1) T93=539=a+92d ----->(2) can u solve these two equations to get a and d ??

hartnn (hartnn):

u can start by subtracting (1) from (2)

OpenStudy (anonymous):

so a=49d-a+92d ?

OpenStudy (anonymous):

*a+49d-a+92d ?

hartnn (hartnn):

yes,thats your right side,where a gets cancelled but whats your left side??

hartnn (hartnn):

and *92d-49d

OpenStudy (anonymous):

oh, so um im not sure

hartnn (hartnn):

ok,subtracting the equations u get 539-238=(92-49)d so \[d=\frac{539-238}{92-49}=??\]

OpenStudy (anonymous):

258?

hartnn (hartnn):

how 258?

OpenStudy (anonymous):

because I got 301/43 = 258?

hartnn (hartnn):

301/43 is actually only 7

OpenStudy (anonymous):

oh worry yeah I got 7

hartnn (hartnn):

so u have d=7 now can put d=7 in any equation and find the value of a

hartnn (hartnn):

238=a+49*7 a=343-238=??

OpenStudy (anonymous):

105?

hartnn (hartnn):

yup :) so u have a=105,d=7 and your general term will be : \[T_n=105+7(n-1)\]

hartnn (hartnn):

yup :) so u have a=105,d=7 and your general term will be : \[T_n=105+7(n-1)\]

OpenStudy (anonymous):

no the answer should be tn=7n-112

OpenStudy (anonymous):

but i dont know how to get that

hartnn (hartnn):

oh,sorry a would be -105 so \[T_n=-105+7(n-1)=-105+7n-7=7n-112\] ok ?

OpenStudy (anonymous):

thank you :)